OpenStudy (anonymous):

The population (in millions) of a certain country can be approximated by the function P(x)=50(1.03)^x , where x is the number of years since 2000. In which year will the population reach 100 million? Hint: An answer such as 2002.4 would represent the year 2002.

OpenStudy (wolf1728):

I think more explanation is needed. The equation (for zero years) generates the number 50. So, the population in the year 2000 is 50 million?

OpenStudy (wolf1728):

P(x)=50(1.03)^x When population = 100 million formula is: P(x)=100(1.03)^x where P(x) would equal 100 100 / 50 = (1.03)^x Taking logs of both sides: log (100) -log(50) = x * log (1.03) 2 - 1.6989700043 = x * 0.0128372247 0.3010299957 / 0.0128372247 = x = 23.4497722504 years So, 23.4497722504 years after the year 2000, the population will be 100 million. So the answer is the year 2023 around the middle of May.